ATLAS UK Cloud Support



Tim Adye (Science and Technology Facilities Council STFC (GB)), James William Walder (Science and Technology Facilities Council STFC (GB))

Password protected (same as (new) OPs Mtg)

ATLAS UK Cloud Support
Zoom Meeting ID
James William Walder
Useful links
Join via phone
Zoom URL

● Outstanding tickets

  • 155856 TEAM atlas RAL-LCG2 less urgent NGI_UK in progress 2022-02-03 09:50:00 RAL-LCG2: deletion errors EGI

    • RAL disk is full, large numbers of deletions of small files.
    • DDM to try and launch larger file deletions.
    • Generally just very busy …
  • 155430 TEAM atlas UKI-SCOTGRID-ECDF less urgent NGI_UK in progress 2022-01-13 14:27:00 UKI-SCOTGRID-ECDF transfer and deletion errors EGI

    • Awaiting large file list to arrive, do be declared as lost
  • 155141 TEAM atlas UKI-LT2-Brunel less urgent NGI_UK in progress 2022-02-01 18:23:00 Transfers from UKI-LT2-Brunel fail with “Internal Server Error” EGI

    • Needing the ok from Brunel that they’ve cleaned the files from their namespace
  • 154806 TEAM atlas UKI-LT2-QMUL less urgent NGI_UK in progress 2022-01-28 09:16:00 UKI-LT2-QMUL SOURCE transfer failures EGI

    • SSL errors remain large source of transfer failures
    • Restarting of some disk server needed
    • Storm added some dev version to repo, causing SRR issues, now reverted
    • Additional storage being added to the site
  • 154543 TEAM atlas UKI-SCOTGRID-ECDF urgent NGI_UK in progress 2021-12-08 12:35:00 DPM storage ACL configuration EGI

    • Still needing an update
  • 154436 TEAM atlas RAL-LCG2 very urgent NGI_UK in progress 2022-02-03 08:15:00 RAL Echo Davs developments EGI

    • More file transfer failures reported due to known multihop limitations
  • 153367 TEAM atlas RAL-LCG2 urgent NGI_UK on hold 2021-12-01 15:37:00 HTTPS on RAL CTA EGI

    • Disucssion on TAPE and Disk tests, following the recommendations of data challenges.


    • RAL

      • Reduced capacity for patching camapaign finished; ATLAS running well, so far …
    • Northgrid

      • NTR
    • London

      • QMUL returning after a few issues.
        • Dark data reported
        • Storm dev version was in repo, also doing the SRR?
    • SouthGrid

      • Xcache issues, followed by power glitch on weekend
    • Scotgrid

      • Xrootd service needed restarting. 503 rucio errors masked this from quickly being spotted

● Ongoing Items

  • TPC with http

    • ATLAS wanting to drop gridFTP by March:
      • Castor <-> Echo would still be ok (if not moved to CTA by then)
      • Echo writes remain on gridFTP (predominantly);
        • With additonal hardware, should be able to switch once installed
      • Glasgow; gridFTP; move to cephfs, or, use patched davs version from RAL
        • Could consider to set up additional SE, and move data across managed by DDM
        • JW to liaise with DDM
  • Storageless Site test (Oxford)

    • NTR (discussed in Storage Mtg)
      • JW to get avaialble plots together by end of Feb
  • LANCS Storage migration

    • Viable endpoint for functional tests should be avaialble by today
    • JW to make final updates to CRIC and run checks before informing DDM



● News round-table

  • Alessandra

    • NTR
  • Dan

    • NTR (discuessed above )
    • Alessandra helping with GPU queue setting
  • Gerard

    • NTR
  • Matt

    • Discussed endpoint avaialbility
  • Sam

    • NTR
  • Stephen

    • NTR
  • Vip

    • NTR



There are minutes attached to this event. Show them.